OpenStudy (anonymous):

Tracy ran for five-sixths hours and Reina ran for one-half hours. How many more hours did Tracy run?

OpenStudy (jdoe0001):

\(\bf Reina\implies 1\cfrac{1}{2}\implies \cfrac{1\times 2+1}{2}\implies \cfrac{3}{2}\\ \quad \\ Tracy\implies \cfrac{5}{6}Reina\implies \cfrac{5Reina}{6}\implies \cfrac{5\times \frac{3}{2}}{6}\implies \cfrac{\frac{5\times 3}{2}}{6}\\ \quad \\ \cfrac{\frac{15}{2}}{\frac{6}{1}}\implies \cfrac{15}{2}\times \cfrac{1}{6}\implies \cfrac{15\times 1}{2\times 6}\)